OpenStudy (anonymous):

What is the y-value of sin(x) when x = 30°? A. -1/2 B. 1/2 C. 1 D. -1

OpenStudy (anonymous):



OpenStudy (anonymous):

sin(30)=opposite/hypotenuse=a/2a=1/2

OpenStudy (anonymous):

thank you @GABI_J

OpenStudy (anonymous):

yw